What can you say about the sets A and B if we know that: a) A ∪ B = A? b) A ∩ B = A? c) A − B = A? d) A ∩ B = B ∩ A? e) A − B = B − A?


1
Expert's answer
2021-10-25T03:53:42-0400

(a) Given:"A \cup B=A"

Using the definition of the union, an element of "A \cup B" is an element that is in A or in B.

"A \cup B=\{x \mid x \in A \vee x \in B\}"

Let x be an element of B.

"x \in B"

By the definition of the union, "x" then has to be an element of "A \cup B" .

"\Rightarrow x \in A \cup B"

By "A \cup B=A" , we then know that x also has to be an element of A :

"\Rightarrow x \in A"

Every element of B is thus also an element of A. By the definition of a subset, we then know that B is a subset of A

"B \subseteq A"


(b) Given: "A \cap B=A"

Using the definition of the intersection, an element of "A \cup B" is an element that is in A and in B.

"A \cap B=\{x \mid x \in A \wedge x \in B\}"

Let "x" be an element of A.

"x \in A"

By "A \cap B=A" , "x" then has to be an element of "A \cap B" .

"\Rightarrow x \in A \cap B"

By the definition of the intersection "x" then has to be an element of B as well.

"\Rightarrow x \in B"

Every element of A is thus also an element of B. By the definition of a subset, we then know that A is a subset of B.

"A \subseteq B"


(c) Given: "A-B=A"

Using the definition of the difference, an element of "A-B" is an element that is in A and not in B.

"A-B=\{x \mid x \in A \wedge x \notin B\}"

Let "x" be an element of B.

"x \in B"

By the definition of the difference, "x" then is not an element of "A-B" ,

"\Rightarrow x \notin A-B"

By "A-B=A" , we then know that "x" also cannot be an element of A :

"\Rightarrow x \notin A"

Every element of B is thus also an element of the complement of A. By the definition of a subset, we then know that B is a subset of "\bar{A}" .

"B \subseteq \bar{A}"

Since the intersection of A and "\bar{A}" is empty and "B \subseteq \bar{A}" , the intersection of A and B then also has to be empty.

"A \cap B=\emptyset"


(d) Commutative law for set identities:

"A \cap B=B \cap A"

The given statement is the commutative law and is thus true for all sets A and B. This then means that we cannot say anything about the sets (since there are no conditions in using the commutative law).


(e) Given: "A-B=B-A"

Using the definition of the difference, an element of "A-B" is an element that is in A and not in B.

"A-B=\{x \mid x \in A \wedge x \notin B\}"

Using the definition of the difference, an element of "B-A" is an element that is in B and not in A.

"B-A=\{x \mid x \in B \wedge x \notin A\}"

Let us assume "x \in A-B" . Then "x \in A" is true and "x \notin A" is false. By the definition of "B-A," "x" can then not be in "B-A" . Since none of the elements of "A-B" are in "B-A" and since "A-B=B-A" , the differences then have to be the empty set.

"A-B=B-A=\emptyset"

The difference A-B does not contain any elements, if all elements of A are also an element of B.

"A \subseteq B"

The difference B-A does not contain any elements, if all elements of B are also an element of A.

"B \subseteq A"

Since "A \subseteq B" and "B \subseteq A" , the two sets then have to be equal.

"A=B"